OpenStudy (anonymous):

An asteroid is traveling in a straight line at a constant speed of 327 m/s. How long will it take for a displacement of 732 km to occur?

OpenStudy (anonymous):

@bennyboy489

OpenStudy (anonymous):

easy bro :) 327ms-1 = 0.327km/s-1 t = d/v t = 732km / .327km/s t = 2238.52 seconds t = 37.30 minutes you could also do this by converting the distance into meters, instead of the velocity

OpenStudy (anonymous):

thank you so much!!! now i get itt :D